OpenStudy (anonymous):

log(2) 2/7 apprioximate value

OpenStudy (solomonzelman):

\(\normalsize\color{red}{ log_2 \frac{2}{7} }\) like this ?

OpenStudy (solomonzelman):

\(\normalsize\color{red}{ \log_2 \frac{2}{7} ~~=}\) \(\normalsize\color{red}{ \log_2 (2 \div 7)~~=}\) \(\normalsize\color{red}{ \log_2 (2)~-~log_2(7)~~=}\) \(\normalsize\color{red}{ 1~-~log_2(7)~~=}\) \(\normalsize\color{red}{ 1~-~2.81~~=}\) \(\normalsize\color{red}{-1.81~~=}\)

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\normalsize\color{blue }{ -1.81 }\)

OpenStudy (anonymous):

log(4) 8+log(4) 2 and simply if possible thats my next one

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \log_48 +\log_42 }\) \(\normalsize\color{blue}{ \log_4(8 \times 2) }\) \(\normalsize\color{blue}{ \log_4(16) }\) \(\normalsize\color{blue}{ \log_4~4^2 }\) \(\normalsize\color{blue}{ 2\times \log_44 }\) \(\normalsize\color{blue}{ 2\times 1 }\) \(\normalsize\color{blue}{ 2 }\)

OpenStudy (anonymous):

ok log(3)x+log(3)-4-log-3 2